OpenStudy (anonymous):

I don't understand how to work this problem out. 5x-7+5 (x+1)=-2x+7

OpenStudy (solomonzelman):

Is it just \(\normalsize\color{blue}{5x-7+5 (x+1)=-2x+7\LARGE\color{white}{ \rm │ }}\) ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

its a linear equation problem @solomonzelman

OpenStudy (nikato):

Do you know how to distribute 5(x+1) ?

OpenStudy (anonymous):

So first combine like terms 5x−7+5(x+1)=−2x+7 5x-7+5x+5 10x-2 then its 10x-2=-2x+7 +2 +2 10x=-2x+9 +2x +2x 12x=9 what is 12 divided by 9

OpenStudy (anonymous):

x=12/5. you can simplify that

OpenStudy (anonymous):

i mean 9

OpenStudy (nikato):

* it's 9 divided by 12 So x=9/12

OpenStudy (anonymous):

Wow nice catch @nikato, my bad.

OpenStudy (anonymous):

thanks @nikato

OpenStudy (nikato):

Mhm.. You're welcome!!